OpenStudy (anonymous):

Two friends are painting a living room. Ken can paint it in 6 hours working alone. If Barbie works alone, it will take her 8 hours. How long will it take working together? Round to the nearest tenth.

OpenStudy (anonymous):

i hated these problems when i was a kid :D but let me try

OpenStudy (amistre64):

the idea is to determine how much work can be done in a given unit of time; in this case, say 1 hour

OpenStudy (anonymous):

never mind :)

OpenStudy (anonymous):

Two words: Harmonic mean

OpenStudy (amistre64):

if K works at a steady rate, then in 1 hour, 1/6 of the room is painted if B works at a steady rate, then in 1 hour, 1/8 of the room is painted togehter in an hour they can paint 1/6+1/8 of the room

OpenStudy (anonymous):

Ken paints 1 room in 6 hours. 1/6 Barbie pains 1 room in 8 hours 1/8 1/6+1/8 = 8/48+6/48 =14/48 So they can pain 14 rooms in 48 hours so 48/14=3.428 hours. Not sure if it's right.

OpenStudy (amistre64):

6+8 ---- is the work per hour ; if we flip that we get hours per work 6x8

OpenStudy (lgbasallote):

1/6 + 1/8 = 1/t multiply everything by t... t/6 + t/8 = 1 8t + 6t = 48 14t = 48 t = 24/7 t = 3.4 hours

OpenStudy (amistre64):

two words harmonic mean is actually 4 words plus some to explain it :)

OpenStudy (anonymous):

Lol, okay you also need to multiply that by \(\frac12\). This is a full answer!